OpenStudy (anonymous):

what is the integral from 0 to infinity of (x^3)/((x^8)+1) help plzzzz

OpenStudy (anonymous):

Note that \(\large\displaystyle \int_0^{\infty} \frac{x^3}{x^8+1}\,dx = \int_0^{\infty}\frac{x^3}{(x^4)^2+1}\,dx\). Now make the substitution \(\large u =x^4\). Can you take things from here? :-)

OpenStudy (anonymous):

but the numerator will u^-1 ?

OpenStudy (anonymous):

No, it will not. If \(\large \color{blue}{u=x^4}\), then \(\large \,du =4x^3\,dx\implies \color{red}{\dfrac{du}{4}=x^3\,dx}\). Thus, \[\large \int_0^{\infty}\frac{\color{red}{x^3}}{(\color{blue}{x^4})^2+1}\color{red}{\,dx} \xrightarrow{u=x^4}{}\int_0^{\infty} \frac{\color{red}{1}}{\color{red}{4}(\color{blue}{u}^2+1)}\color{red}{\,du} = \frac{1}{4}\int_0^{\infty}\frac{1}{u^2+1}\,du\] Does this clarify things? Do you think you can take things from here? :-)

OpenStudy (anonymous):

yeah i thought about that when i differentiate but when will subtitute back to x^4 ?

OpenStudy (anonymous):

after i take the limit

OpenStudy (anonymous):

You can change your limits of integration so you don't have to back substitute after finding the integral. In this case, when you change the limits of integration, you get the same limits back.

OpenStudy (anonymous):

is it Arctan of U around 0 and infinity

OpenStudy (anonymous):

In particular, if \(x=0\), then \(u(0) = 0^4 = 0\); thus the lower limit for the variable u is 0. Likewise, as \(x\to infty\), we see that \(\displaystyle \lim_{x\to\infty} u(x) = \lim_{x\to\infty} x^4 = \infty \); hence the upper limit for the variable u is \(\infty\). And yes, \(\large\displaystyle \int_0^{\infty} \frac{1}{u^2+1}\,du = \left.\bigg[\arctan u\right]\bigg|_{0}^{\infty}\).

OpenStudy (anonymous):

whats my new limit of integration plzz

OpenStudy (anonymous):

nice so it wont change

OpenStudy (anonymous):

Exactly; that's the nice thing about changing the limits of integration in this case.

OpenStudy (anonymous):

However, don't forget the factor of \(\large \dfrac{1}{4}\) in your integral!

OpenStudy (anonymous):

you are the besssssst

OpenStudy (anonymous):

I just want to make sure you got the right answer; what did you get as the final answer? :-)

OpenStudy (anonymous):

yesir its 1/4 times pi/2

OpenStudy (anonymous):

pi/8

OpenStudy (anonymous):

Exactly, which is \(\large\dfrac{\pi}{8}\). I hope this made sense! :-)

OpenStudy (anonymous):

yes it did u are truly my savior